What are the length, width, and height of the prism if the volume is 3x3 + 9x2 + 6x?
Answer by ewatrrr(24785)

3x^3 + 9x^2 + 6x = 3x(x^2 + 3x + 2) = 3x(x+1)(x+2)= Lwh